A solution of sodium hydroxide has a concentration of 60 g dm-3 . What is the molarity of the alkali?

[ Relative atomis mass : H=1, O=16, Na=23]


1
Expert's answer
2021-08-06T03:44:02-0400

m= 60 g

M(NaOH)=23+16+1= 40 g/mol

"n = \frac{m}{M} \\

n=\frac{60}{40}=1.5 \;mol"

Molarity = 1.5 mol/L = 1.5 M

Answer: 1.5 M